Approximately 100 g of PbCl4 reacted excess NH4Cl according to the following equation: PbCl4 + 2 NH4Cl → (NH4)2PbCl6. If the yield was 87.0%, how many grams of product were obtained?



please include calculations

1
Expert's answer
2022-02-01T19:24:02-0500

The molar masses are:

PbCl4 - 349.01 g/mol;

(NH4)2PbCl6 - 455.99 g/mol.


"m \ (NH_4)_2PbCl_6=100\ g\ PbCl_4\times\frac{1\ mol\ PbCl_4}{349.01\ g\ PbCl_4}\times\frac{1\ mol\ (NH_4)_2PbCl_6}{1\ mol\ PbCl_4}\times\frac{455.99\ g\ (NH_4)_2PbCl_6}{1\ mol\ (NH_4)_2PbCl_6}\times0.870=114\ g"


Answer: 114 g